OpenStudy (anonymous):

please help! Find all solutions to the equation (sin x–cos x)²=3

OpenStudy (anonymous):

\[(a-b)^2=a^2-2ab+b^2\]Apply this. ----------------------------------

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

You can let a=sin(x) and b=cos(x), try. :) then, substitute after. tell me, what did you get.

OpenStudy (anonymous):

\[\large (\sin x -\cos x)^2-3=0\] \[\large \sin^2 x-2\sin x \cos x+\cos^2 x-3=0\] \[\large \sin^2 x-2\sin x\cos x+1-\sin^2 x-3=0\] \[\large -2\sin x\cos x-2=0\] \[\large 2\sin x\cos x=-2\] Using your trig identities: \[\large \sin 2x=2\sin x\cos x\] Therefore: \[\large \sin 2x=-2\] \[\large 2x= no solution\]

OpenStudy (tkhunny):

We really should spend more time on this transformation. It is quite useful... \(\sin(x) - \cos(x) = \sqrt{2}\cdot \cos\left(\dfrac{3\pi}{4} - x\right)\)

OpenStudy (anonymous):

That's where I was gonna follow up to say use auxillary angles.

OpenStudy (anonymous):

ANd you got it wrong way around inside the cosine.

OpenStudy (anonymous):

thank you all so much, i will go over this now. i dont know who to give best answer wish i could give it to you both

OpenStudy (anonymous):

From @Azteck \[\sin(2x)=-2\]\[2x=\sin^{-1}(-2)\]\[x=\frac{ \sin^{-1} (-2) }{ 2 }\] Inverse function of sine is \[-1\le x \le1\]So, \[\sin^{-1} (-2)=undefined\]OR it's not in the domain

OpenStudy (anonymous):

\[\sin x -\cos x=\sqrt{2}\sin(\frac{x-\pi}{4}=\sqrt{3}\]

OpenStudy (anonymous):

Wth... computer error. Sorry wrote it wrong.

OpenStudy (tkhunny):

@Azteck No, that is correct. There may be other equivalent forms.

OpenStudy (anonymous):

\[\sin x -\cos x=\sqrt{2}\sin(x-\frac{\pi}{4})\]

OpenStudy (anonymous):

I never seen that before @tkhunny The other option with cos I could think of would be the same as what I wrote just then but with a + sign and sin becomes a cos.

OpenStudy (anonymous):

And there will be no solution as well.

OpenStudy (tkhunny):

Equivalent forms are equivalent. They don't care how you find them or if anyone has met them, before.

OpenStudy (anonymous):

so the answers are no solution and sinx−cosx=2√sin(x−π4) ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

That was another method of finding solutions to your question.

OpenStudy (anonymous):

oh ok, so what are the final answers then

OpenStudy (anonymous):

@tkhunny They care about how you find them. That's called working out.

OpenStudy (anonymous):

And if you say it's correct and equivalent, then good for you, well done.

OpenStudy (tkhunny):

@Azteck Meh. Humans care. Forms and Solutions don't care. :-)

OpenStudy (anonymous):

Lol. True!

OpenStudy (anonymous):

sorry im still confused, what are the answers in that case?

OpenStudy (tkhunny):

In all cases, no matter how you work it, the RANGE of \((\sin(x) - \cos(x))\) is \([-\sqrt{2},\sqrt{2}]\). Now that we know that, we can see that the RANGE of \((\sin(x) - \cos(x))^{2}\) is \([0,2]\). A little thought on this result should suggest that it will never take on the value 3. Done. No solution.

OpenStudy (anonymous):

thanks so much :)